OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-1, -4) and (4, 11)

OpenStudy (anonymous):

Slope is 3.

OpenStudy (anonymous):

11 - (-4)/ 4 - (-1) = 15/5 = 3

OpenStudy (anonymous):

Now pick either point: 11 = 3(4) + b 11 = 12 + b -1 = b So, y = 3x -1 is your answer.

OpenStudy (anonymous):

Note: I used y = mx + b, and plug in 11 for y, 3 for m and 4 for x.

OpenStudy (anonymous):

thank you